A population consists of the numbers 2 4 8 10 and 5.Let us list all the possible samples of size 3 from this population and construct the sampling distribution of the sample mean


A.How many possible samples can be drawn


B.Construct the sampling distribution of sample means


C.Make a probability distribution of the sample mean (Sampling distribution of the sample means)

1
Expert's answer
2022-03-10T20:43:02-0500

A - There are 4 possible samples.

B - (4+8+10+5)/4 =6.75

C - μ = Mean. σ = Standard Distribution.

6.75*5 = 33.75